Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day05


Crew 99 Log Book for February 2, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting


Crew Physical Status: Good


Time Departed/Returned from EVA: EVA #6 12:00-13:00; EVA #7 14:30-16:30


Brief Narrative of Field Mission Results: EVA planner test; geological

measurements and sample site investigation


EVA Data/Interpretations: as per EVA reports


Engineering/Hab Maintenance: as per engineering report; problems with

water line/greywater seem to have been resolved


Report Transmission Schedule: Engineering, EVA #6, EVA #7, Chef's,

Commander's. By 8:30pm


Plans for Tomorrow: to be decided


Inventory: n/a


Support Requested: none


Miscellaneous: Happy Groundhog Day!!




Commander's Report

Donna Viola Reporting

After a long night of continuing to fiddle with out primary LAMBDA experiment, the team slept in a little this morning. However, by 10am, we were all set to get started with our first experiment to test for microbial metabolism in a soil sample we collected approximately a kilometer south of the Hab.

After everything was set up and running, Andie and Max headed off for the Striped Dragon for another test of the EVA mission planner. Back at the Hab, there were some problems with freezing pipes due to the cold weather, but that seems to have been resolved (here on Mars, we're a little too far from civilization to know whether everyone's favorite groundhog saw his shadow this morning, but from what we can tell, it seems like winter is going to stick around for a little while longer).

This afternoon, Heidi, Kevin, and I set off on the ATVs to collect some geological data and to explore some potential future sample sites - and returned to some nice mugs of hot chocolate!

Dinner was prepared by Max and Sukrit, our crew astronomers - who are hoping to get some observatory time in tonight.


Engineering Report

Kevin Newman Reporting

Generator/Electricity:

·          Kitty  (Diesel Generator) (Used? Yes/No): Yes

·          Honey   (Gas Generator) (Used? Yes/No): No

·          Battery Charge Level:

o    10.59 V everstart

o    12.37 V armor plate

o    13.32 V green

Notes/Comments (include how many generator runs, approximately when,
and any times you needed to turn the generator on/off):
1 generator run, all is well

There was a small leak in the pressurized walkway, nearly lost an
engineer, crisis averted.

Fuel Status:

·          Diesel (Full/0.75/0.5/0.25/Empty):2/7

·          Propane Tank (Full/0.75/0.5/0.25/Empty):76%

·           Gasoline Tank (Full/0.75/0.5/0.25/Empty): Horizontal
raised drum EMPTY. Three 5-gallon containers full


 Notes/Comments:

We are running low on oxygen. Please advise.

Oil:

·          Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

·          Honey Oil Quantity (# of Quarts in storage): Leave it to DG

·          ATV Oil Quantity (# of Quarts in storage): 6 quarts on the
engineering work bench

·          Car Oil Quantity (# of Quarts in storage):

Water Status:

·          Outside Potable Water Tank Level (inches from bottom): 13

·           Trailer Potable Water Tank Level (inches from bottom): 1/5

·           Hab Potable Water Tank Level (inches from bottom): 12

·           Potable Water Meter Reading: 415891

Notes/Comments:
Pump from outside potable water tank to hab potable water tank not
operational this morning 9:00, likely frozen. Attempted again around
17:00 and working properly again.

GreenHab:

·          Condition of plants in Tank 1: Nothing

·          Condition of plants in Tank 2: Ok

·          Condition of plants in Tank 3: Good

·          Greenhab notes/ comments:

Observatory:

·          Telescope Used? (Yes/no) No

 Observatory Notes:

Transportation:

Hab Car (NOT RENTAL)

·          Used (yes/no):Yes

·          Oil Checked (yes/no): No

·          Tire Status: Not checked

·          Odometer at end of day: 123597.9

·          Notes/Comments on Hab Car:

o    Started and driven briefly for the sake of keeping the engine
warm in cold weather.

ATV 1 (Opportunity):

·          Used: Yes

·          Oil Checked: Yes

·          Fuel Consumed: Replaced

·           Tire Status: Checked, Filled

·           Comments:

 ATV 2 (Spirit):

·          Used: Yes

·          Oil Checked: Yes

·          Fuel Consumed: Replaced

·          Tire Status: Checked, Filled

·          Comments:

ATV 3 (Viking I):

·          Used: No

·          Oil Checked: No

·          Fuel Consumed: None

·          Tire Status: Not Checked

·          Comments:

ATV 4 (Viking II):

·          Used: Yes

·          Oil Checked: Yes

·          Fuel Consumed: Replaced

·          Tire Status: Checked, Filled

·          Comments:

Heating and Ventilation:

·          Thermostat setting upstairs (Farenheit): 70

·          Actual temperature upstairs (Farenheit): 75

·          Thermostat setting downstairs (Farenheit): 65

·          Actual temperature downstairs (Farenheit): 62

Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All good

General Engineering
1. Notes/Comments: Its still really cold!!!

Clear Skies!



EVA 6 Report

Andie Gilkey Reporting


Time: 12:00-12:54


Crew Members: Andie Gilkey, Max Fagin

Site Location:  UTM: 518250 4250100 and UTM: : 518333 4250900
 Transit Mode: foot

Objectives: Measure how accurately the mission planner predicts astronaut energy consumption, distance traveled, and EVA time for two different locations.

Results: We were able to successfully measure astronaut heart rate throughout the length of the missions, as well as distance traveled and EVA time for both locations.

Lessons Learned: Bring a pencil when you have to record data on your EVA...pens may run out of ink.


EVA 7 Report

Heidi Beemer Reporting

EVA # 7

Time: 14:30-6:35

Crew Members: Heidi Beemer, Donna Viola, Kevin Newman

Site Location:  Lowell Highway and surrounding roads

Objectives: Measure joint separation at geological location from yesterday. Locate Cactus Road and conduct an exploratory recon of surrounding roads and junctions to the NW of Hab.

Results: Joints were measured. Results to come.  We missed the junction to cactus road but with some land navigation skills and back tracking we were able to locate the junction between Brahe Highway and Lowell Highway (N38.44184 W110.79605). The junction between unnamed road and Brahe Highway was also found (N38.43909, W110.81619) We were also able to locate Sagan Street and other roads to the NW.

Lessons Learned: It is bitter to receive an email from about not going on a long EVA when it is cold and windy right after coming back from a long, cold, windy, Eva.

Request Support:  Is there already a data base of GPS points of road junctions. If yes can we get this. If no would this be worth us compiling?



Journalist Report

Sukrit Ranjan Reporting

Today, Crew 99 started it's primary experiment: using microbial fuel cells (MFCs) to look for life!

After rapid Macguyvering by our engineering team to resolve an agitation problem, we put our soil samples in the MFC and are now tracking voltage as a function of time.

Unfortunately, our fix requires us to manually turn on the agitation mechanism -- once ever hour. All night long. We're taking shifts.

In addition to further tests on our mission planner software and geological and exploratory EVAs, Crew 99's astronomers made progress in restoring Musk observatory. Now, we just need to figure out why our hand controller is talking in Greek...


Cook's Report

Max Fagin Reporting

Max Fagin
2/02/11

1a) Today was
-A cooking day.

1b) Was today a special day celebrated at the main meal with special
food or activity?

- No

2)  List the foods served at today's main meal (usually the evening
meal), giving full names of each.

- Pre-cooked Bacon, 30 pieces.  Not heart attacks to report yet.
- Chicken and Vegetable Soup


3) List any main meal foods not finished by the crew (leftovers stored
for later use or discarded)  (Example:  2 servings potato soup stored;
2 servings rice burned/inedible, discarded)

-Rice sweet-cakes were prepared for after dinner dessert, but were far to undercooked to be edible.  1 cup of undercooked rice was left unconsumed and disposed of after dinner.

4)  Recipes for anything more complicated than following package
directions exactly, or rehydrating.

- Rice sweet cakes:  Cook rice, add brown sugar, vanilla extract and powdered sugar.  Bake at 350 °F until crispy golden brown and serve with syrup.

5)  Comments/questions for Kim and Jean, the food study investigators

-None q